# A Disproof of Large Language Model Consciousness: The Necessity of Continual Learning for Consciousness

### v1

@@ A Disproof of Large Language Model Consciousness: The Necessity of Continual Learning for Consciousness
Erik Hoel hoelerik@gmail.com Bicameral Labs, MA, USA
@@ 1 Abstract
The requirements for a falsifiable and non-trivial theory of consciousness significantly constrain such theories. Specifically, recent research on the Unfolding Argument and the Substitution Argument has given us formal tools to analyze requirements for a theory of consciousness. I show via a new Proximity Argument that these requirements especially constrain the potential consciousness of contemporary Large Language Models (LLMs) because of their proximity to systems that are equivalent to LLMs in terms of input/output function; yet, for these functionally equivalent systems, there cannot be any non-trivial theory of consciousness that judges them conscious. This forms the basis of a disproof of contemporary LLM consciousness. I then show a positive result, which is that theories of consciousness based on (or requiring) continual learning do satisfy the stringent formal constraints for a theory of consciousness in humans. Intriguingly, this work supports a hypothesis: If continual learning is linked to consciousness in humans, the current limitations of LLMs (which do not continually learn) are intimately tied to their lack of consciousness.
2 A Falsifiability-First Approach
@@
In the past decade, consciousness research has taken a “structural turn” and become focused on mathematics and formalisms [kleiner2020mathematical, kleiner2024towards]. Much of this has to do with the influence of Integrated Information Theory (IIT) [albantakis2023integrated] becoming a popular and mathematically well-formalized theory of consciousness [barbosa2020measure, kleiner2021mathematical]. In turn, many of the criticisms of IIT have depended on its detailed particulars and formalisms and whether it is scientifically testable [bayne2018axiomatic, fleming2023integrated]. This has led to increased interest in defining things like minimal “toy models” of consciousness in order to differentiate potential theories and test their boundaries [hoel2024world, albantakis2025utility], and identify what sort of testing requirements might constrain theories of consciousness.
@@
This leads to an interesting question: Can a theory of consciousness be so strongly constrained by requirements for falsification and testability that its overall nature can be deduced?
@@
Increasingly, this seems a necessity. Right now, the field of consciousness research has hundreds of existing theories [seth2022theories, kuhn2024landscape], and no agreed-upon way to empirically distinguish them. The search for the neural correlates of consciousness [crick1990towards, koch2016neural] has been shown to have a serious problem wherein even tightly-controlled and well-funded adversarial collaborations fail [cogitate2025adversarial], such as the recent comparison of the highly-popular theories IIT and Global Neuronal Workspace Theory [mashour2020conscious, baars1993cognitive]. Instead, that adversarial collaboration led to accusations of pseudoscience [fleming2023integrated, tononi2025consciousness]. If consciousness research is to make serious progress in our lifetime, it must winnow the field via constraints (for a list of some such constraints, see [doerig2021hard]).
@@
Working backwards from “What kinds of theories of consciousness are falsifiable and non-trivial?” is a far less explored option for deriving a theory (or class of theories). In this, scientists would act like artists drawing the negative space around consciousness to see its outline clearly, with minimal metaphysical assumptions and no adherence to any particular theory.
Another reason to pursue this falsification-first approach is that we want questions about consciousness that the contemporary state of consciousness research cannot answer yet. For instance, the question whether or not contemporary LLMs are conscious has become critical, as one of the major risks of the current uncertainty around AI and consciousness is falsely attributing contemporary LLMs (like Claude or ChatGPT) moral status (which may also provoke increased incidents of AI psychosis) [seth2024conscious, morrin2025delusions]; alternatively, they may themselves be moral patients [sebo2025moral]. If it is possible to prove that there is no theory of consciousness that is non-trivial and falsifiable that could exist for LLMs, that is much faster progress than trying to assess the probability of consciousness by assigning assumption-dependent probabilities, or applying the many (often contradictory) theories that do currently exist [butlin2023consciousness], as those are not just unproven, but even worse, they are empirically indistinguishable in practice.
@@
A road to a potential disproof of LLM consciousness begins with one particularly influential criticism of IIT: The “Unfolding Argument” by Doerig et al. [doerig2019unfolding] has led to an evocative debate in the literature on precisely how to falsify theories of consciousness. Specifically, the argument is based on the universal approximation theorem [hornik1989multilayer], and argues that any given Recurrent Neural Network (RNN) can be “unfolded” (or just replaced) via some functionally equivalent (in terms of input/output) Feedforward Neural Network (FNN). This presents a significant problem for theories like IIT, since an RNN might have a certain degree of integrated information, yet its input/output “twin” FNN would have zero integrated information.
One might argue that the FNN’s equivalence is only input/output, and that a theory’s predictions about consciousness could still distinguish between the two cases by making differing predictions based on their internal workings. Unfortunately, this is equivalent to a theory assuming what it seeks to prove, since it ends up being equivalent to judging a theory of consciousness on the basis of internal brain (or neural network) activity alone and rejecting cases where behavior (or report, or more generally output) would falsify the theory.
@@
However, the Unfolding Argument alone only gives us a hint that internal “causal structure” theories face significant challenges. In later work by Johannes Kleiner and myself [kleiner2021falsification], we pointed out two further conclusions: First, that unfolding being possible was not some specific trick based on that theorem, but a consequence of the independence between what a theory uses to make predictions about consciousness (like the integrated information of IIT) and how a scientist infers the actual states of conscious experiences during experiments (via things like report or behavior or more broadly, outputs). We argued that this made it apply to a broader swath of theories than initially thought, and that we could think of the issue as a framework wherein alternative systems were “substituted” in during empirical testing. Additionally, we pointed out that there is also a risk on the other side of having too much dependency between a theory’s predictions and a scientist’s inferences, which would create theories that are unfalsifiable. In other words, there are two “horns” to the problem of falsifiability for theories of consciousness.
@@
The dual-horned “Kleiner-Hoel dilemma” established in [kleiner2021falsification] goes like this: A successful theory of consciousness must not be a priori falsified by the class of arguments like the Unfolding Argument (broadly, any such “universal substitution,” wherein unfolding is just one such possible substitution). However, the same theory must also not be unfalsifiable in principle. This creates only a small space through which to pass; theories of consciousness are “caught between Scylla and Charybdis.”
@@
On one side, theories like IIT [tononi2012integrated] or recurrent processing theories [lamme2006towards] are a priori falsified by substitutions like unfolded neural networks (since pretty much all the predictions of a theory would change under such a substitution, while the experimenter’s inferences based on reports or behavior would stay exactly the same, creating universal mismatches between predictions and inferences and thus resulting in falsification). Meanwhile, theories like behaviorism or global functionalism (both being based solely on input/output mattering for consciousness) are unfalsifiable, because predictions of the theories and inferences about consciousness from experiments can never differ (e.g., if your theory of consciousness is that consciousness is identical to consciousness-like behavior, how could it ever be falsified?).
@@
This was a strong challenge to the consciousness literature that was never addressed. The primary reply against it has been simply that the Kleiner-Hoel dilemma is too strong and therefore cannot possibly be true, at least without ruling out consciousness research in general [usher2023philosophical]. This is a debatable reason to reject a conclusion, but more importantly, it falls by the wayside once an appropriate theory is introduced (even the broadest sketch of one) that can navigate the Kleiner-Hoel dilemma. Critically, if the Kleiner-Hoel dilemma is navigated successfully in humans, but provably cannot be navigated in LLMs (due to their close resemblance to systems that we know cannot possess consciousness except in trivial ways), this would form a strong argument against LLM consciousness. In other words, there may be properties of human brains that allow them to “pass through” the Kleiner-Hoel dilemma unscathed, but LLMs, lacking those properties, cannot.
That is what this paper shows. In the process, it constrains our theories of consciousness significantly, and gives us a disproof of modern LLM consciousness as well.
@@
First, I overview the formal framework for testing theories of consciousness introduced in [kleiner2021falsification], and demonstrate how the Kleiner-Hoel dilemma naturally arises from it. Next, the strength of reasoning in a falsification-first way for theories of consciousness is demonstrated. I show that the Kleiner-Hoel dilemma has extra argumentative force in LLMs, because LLMs are much closer than human brains in “substitution space” to things like lookup tables that non-trivial theories of consciousness can’t possibly judge as conscious. That is, there is no theory of consciousness available for contemporary LLMs that would not fall on one of the horns of the Kleiner-Hoel dilemma. Therefore, there is no scientific theory of consciousness which could judge contemporary LLMs as conscious. Therefore, they are not conscious.
@@
The strength of this argument is that it is entirely agnostic to any assumptions about particular theories of consciousness being true, nor does it make assumptions about the “specialness” of biology, or the possibility of AI consciousness in general, and has few, if any, philosophical commitments (other than that theories of consciousness should be non-trivial and testable).
@@
Additionally, this paper shows that the Kleiner-Hoel dilemma has a potential solution in humans by grounding a theory of consciousness not in causal structure like IIT, nor functional or dynamical features like Global Workspace Theory, but in continual learning. This fits well with the recent observation by O’Reilly-Shah et al. that a static FNN cannot act as a universal substitution for a plastic RNN [o2025caveat]. However, the conclusions from those recent proofs only show that plastic RNNs are protected from the specifics of the Unfolding Argument—they do not fully distinguish whether the recurrence or plasticity is the cause, nor ask if this protection extends to other universal substitutions. Here I show that grounding a theory of consciousness in the broader context of continual learning explicitly avoids both horns of the Kleiner-Hoel dilemma. This is because a theory grounded in learning is a case of “lenient dependency.” The possibility of lenient dependency was speculated about in [kleiner2021falsification], but no concrete possibilities were available at the time.
@@
I show that this result implies that scientific theories of consciousness are, until other cases of lenient dependency are identified, strongly constrained to not merely being tested in systems that can learn (like plastic RNNs), but more generally that we can constrain theories of consciousness to (at minimum) involving continual learning. Such a result significantly winnows the field of consciousness research and illuminates the path toward less popular, but ultimately more scientific, theories of consciousness. It also implies that the vast majority of the search for the neural correlates of consciousness may be incommensurate, because testing input/output behavior alone is not how one empirically tests for consciousness.
@@
Intriguingly, this research indicates that one major reason deployed LLMs lack consciousness is their static structure, and that unlocking continual learning, already a significant target in AI research [kudithipudi2022biological], could possibly change this disproof. Therefore, this paper provides grounds to believe that the current cognitive limitations of LLMs are based on their lack of consciousness, which is connected to their lack of continual learning.
3 A Formal Falsification Framework for Theories of Consciousness
@@
The line of reasoning here will make use of the framework for falsification of theories of consciousness that I developed with Johannes Kleiner [kleiner2021falsification]. This is a formalized mathematical framework for testing theories of consciousness, based on treating theories as mapping a space of possible experiences onto physical predictions. Its advantage is that it is abstract, general, and (relatively) simple.
@@
Here, I leave much of the background logic and reasoning to previous work and deal only with the high-level details that are relevant (see [kleiner2021falsification] for those details). For our purposes what matters is that testing a theory of consciousness is based on comparing two functions: predictions and inferences.
@@
First, what are predictions? Predictions pick out a particular element of the space of possible experiences, $E$. The correspondence (here, for simplicity, referred to as a function) that picks this out is a mapping from experimental observations about the system’s workings, $O$ (such as neuroimaging data), to the space of experiences, $E$; this mapping is dubbed pred, for “prediction.”
@@
However, predictions alone cannot falsify a theory. They need to be compared to inferences about consciousness, either explicit or implicit.
Usually, these inferences are based on verbal report, behavior, or, more recently, even “no-report paradigms” [tsuchiya2015no]. But even “no-report paradigm” cases (based in things like eye movement) rely on correlating with report first. In a psychology lab, button presses to indicate what was flashed in a masking stimulus are inference data; in another context, like in assessing LLM consciousness, the outputs of an artificial neural network would be inference data. Inferences are, essentially, empirical reasons to believe a system might be conscious irrespective of any given theory or any strong proto-theoretical concepts based on some larger class of theories.
Overall, inferences are represented by another function, which maps between data generated by observations about the input/output of a system (like its incoming “sensory” information, and its responsive behavior or reports) and the space of experiences $E$ (delineated by the theory of consciousness being tested):
@@
Figure 1: A theory of consciousness is not falsified (sometimes called “experimentally confirmed,” although that language is often not appropriate) when inference from behavioral reports and prediction from neural data converge on the same experience $E$.
Obviously, this formal setup is an extremely idealized and abstracted version of testing a theory of consciousness (for a defense of the irrelevancy of these details, see [kleiner2021falsification]).
It is important, however, that a falsification framework is not overfitted or over-sensitive. For the purposes relevant to the paper, I do not require that falsification triggers from just any possible mismatch between pred and inf. For instance, imagine a theory of consciousness that performs very well in all other respects and we are starting to believe scientifically. Then, we apply it to some systems for which we cannot access consciousness as well (say, plants). This gives us a (to many) surprising result: that plants are conscious to some degree. However, it may be that no experimental inference from plants can fully support the weight of the claim of the theory, and therefore, there are technically some mismatches. The point here is not to say this should happen with a theory, or is totally non-problematic, but just that, at this time, the falsification framework herein can be optionally permissive about some occasional mismatches.
On the other hand, if, e.g., a theory of consciousness predicts that extremely simple systems have degrees of consciousness greater than that of a human [aaronson2014unconscious, tononi2014conscious], such a radical mismatch that would count as actual falsification—but only in light of the inference that a system’s behavioral or input/output simplicity is evidence against such a high degree of consciousness. Ideally, such inferences should always be made explicit when reasoning about consciousness through the lens of falsification. However, it is important to note that keeping inferences as neutral as possible is important, or else testing a theory of consciousness would lose its desired empirical “universality” [kanai2024toward].
Importantly, it is obvious that if inf is held constant, but pred is varied significantly, then this will create consistent divergence (mismatches) across the entirety (or vast majority) of a theory’s predictions. This is pathological behavior and should necessarily falsify a given theory, as now every prediction is wrong (or the vast majority are wrong) for a system with identical (or extremely similar) inferences. Specifically, in [kleiner2021falsification] we took the approach that a “universal substitution” (of one system for another) in an experimental setup (within sensible allowances), but in which inferences do not change, but all predictions do change, would completely falsify a particular theory. We called such a possibility a “universal substitution” and confined the analyses to such cases in order to avoid more debatable instances. E.g., it is unclear to what degree “partial substitutions” triggering mismatches should count as falsification, since it may make sense to sometimes err on the side of a well-confirmed theory in edge cases.
@@
As we showed, substitutions are guaranteed to exist as long as predictions and inferences are totally independent. Another way of saying this: Just from knowing the structure of Integrated Information Theory, and that the calculation of $\Phi$ is entirely screened off from the input/output, there can be systems wherein $\Phi$ differs but the input/output does not. However, the reason why this appears so prominently in IIT is simply that IIT is formally well-developed. Other theories will eventually become formally developed, and have to specify similar predictions to $\Phi$, and then they will fall victim to the substitution argument.
[…]
Figure 2: The Substitution Argument: Changing the system (and thus significantly changing the prediction function), either by transformation or swapping in a different system, all while holding inference data constant, is always possible as long as inferences and predictions are completely independent. Here, following a substitution, the predicted experience shifts from $E$ to $E^{\prime}$, but the inferred experience remains $E$, signifying the mismatches that falsify the theory.
What are some examples of universal substitutions?
  • Lookup tables that implement the same input/output function of a system directly [searle1980minds].
  • Static FNNs that implement the same input/output (or approximate it precisely) [doerig2019unfolding].
  • Turing machines implementing functionally-identical programs or algorithms that model another system (again, or approximate it very closely) [turing1936computable].
  • Potentially more exotic options, like substitutions based on the AIXI model of universal intelligence [hutter2000theory] or even cases of a brain in quantum superposition (a ”Schrödinger’s Zombie”) [brown2019schrodingers].

### v2

@@ A Disproof of Large Language Model Consciousness: The Necessity of Continual Learning for Consciousness
Erik Hoel erik@bicameral-labs.org Bicameral Labs, MA, USA
@@ 1 Abstract
Scientific theories of consciousness should be falsifiable and non-trivial. Recent research has given us formal tools to analyze these requirements of falsifiability and non-triviality for theories of consciousness. Surprisingly, many contemporary theories of consciousness fail to pass this bar, including theories based on causal structure but also (as I demonstrate) theories based on function. Herein I show these requirements of falsifiability and non-triviality especially constrain the potential consciousness of contemporary Large Language Models (LLMs) because of their proximity to systems that are equivalent to LLMs in terms of input/output function; yet, for these functionally equivalent systems, there cannot be any falsifiable and non-trivial theory of consciousness that judges them conscious. This forms the basis of a disproof of contemporary LLM consciousness. I then show a positive result, which is that theories of consciousness based on (or requiring) continual learning do satisfy the stringent formal constraints for a theory of consciousness in humans. Intriguingly, this work supports a hypothesis: If continual learning is linked to consciousness in humans, the current limitations of LLMs (which do not continually learn) are intimately tied to their lack of consciousness.
2 A Falsifiability Approach to Consciousness
@@
In the past decade, a subfield of consciousness research has taken a “structural turn” and become focused on mathematics and formalisms [48, 49]. Much of this has stemmed from the influence of Integrated Information Theory (IIT) [2] becoming a popular theory of consciousness, since it is expressed in a mathematically formalized way [11, 47]. In turn, many of the criticisms of IIT have depended on its detailed particulars and mathematical formalisms and whether it is scientifically testable [12, 32]. These debates have led to increased interest in defining things like minimal “toy models” of consciousness in order to differentiate potential theories and test their boundaries [38, 4], and identify what sort of requirements might constrain theories of consciousness.
@@
This structural turn gives rise to a historically novel question: Can a theory of consciousness be so strongly constrained by requirements for falsifiability and non-triviality that its overall nature can be deduced?
@@
Increasingly, such an approach seems a necessity. Right now, the field of consciousness research has hundreds of existing theories [65, 54], and no agreed-upon way to empirically distinguish them. The search for the neural correlates of consciousness [27, 51] has been revealed to have a serious problem wherein even tightly-controlled and well-funded adversarial collaborations fail [26], such as the recent comparison of IIT and Global Neuronal Workspace Theory [57, 9], which led to accusations of pseudoscience [32, 67]. If consciousness research is to make serious progress in our lifetime, it must winnow the field via constraints (see, e.g., [30]).
@@
Compared to other approaches, working backwards from “What kinds of theories of consciousness are falsifiable and non-trivial?” is a far less explored option for deriving a class of viable theories. In this, scientists would act like artists drawing the negative space around consciousness to see its outline clearly, with minimal metaphysical assumptions and no adherence to any particular theory.
Another reason to pursue this falsification-first approach is that we want questions about consciousness that the contemporary state of consciousness research cannot answer yet. Most prominently, the question whether or not contemporary LLMs (like ChatGPT or Claude or Gemini) are conscious has become suddenly critical. There are major risks associated with getting this question wrong. The more scientifically conservative perspective is that contemporary LLMs are not conscious and yet may (depending on prompt) seem conscious and so be taken as if conscious. This has a myriad of risks, which include increasing the risk of AI psychosis, overestimation of their capabilities and import, inappropriate practices or regulation, and misleading scientific beliefs about human consciousness [66, 58]. On the other hand, if contemporary LLMs were conscious they could be considered moral patients [64] and be due considerations like conversational “exit rights” [7] as well as other potential rights against mistreatment.
It seems unlikely that empirical data from LLMs alone can answer the question of their consciousness. LLMs are highly prone to confabulation [41]. And while there is a small subfield studying LLM “introspection” [14], the best evidence for this introspection is in the form of statistically uncommon anomaly detection of injected concepts, for which there could be shallow mechanistic explanations [8]. There is not even agreement over whether (low-resolution) memory stored from a forward pass of previous internal processing actually counts as anything like human “introspection,” nor if it should be associated with consciousness.
Therefore, most have assumed that answering the question of LLM consciousness requires a scientific understanding of consciousness beyond the current day (e.g., scientific consensus around a particular theory of consciousness). However, I show that this is not necessary. Instead, it is possible to prove that there is no theory of consciousness that is non-trivial and falsifiable that could exist for LLMs. This represents much faster progress than trying to assign assumption-dependent probabilities of consciousness, or trying to apply some set of the many (often contradictory) theories that do currently exist [20].
@@
This road to a disproof of LLM consciousness begins with one particularly influential criticism of IIT: The “Unfolding Argument” by Doerig et al. [31] has led to an evocative debate in the literature on precisely how to falsify theories of consciousness. Specifically, the argument is based on the universal approximation theorem [39], and argues that any given Recurrent Neural Network (RNN) can be “unfolded” into some functionally equivalent (in terms of input/output) Feedforward Neural Network (FNN). This creates a pathological scientific situation for theories like IIT, since any given RNN might have a certain degree of integrated information, yet its input/output “twin” FNN would have zero integrated information.
@@
However, the Unfolding Argument alone only brings to light that internal “causal structure” theories like IIT face significant challenges. In later work by Johannes Kleiner and myself [45], we reached two further conclusions. First, that unfolding being possible was a consequence of the independence between what a theory uses to make predictions about consciousness (like the integrated information of IIT) and how a scientist infers the actual states of conscious experiences during experiments (via things like report or behavior or more broadly, outputs). We argued that this made the unfolding argument apply to a broader swath of theories than initially thought, and that the best conceptualization of the issue is a framework wherein alternative systems can be “substituted” in during empirical testing in order to reveal pathologies around falsifiability. Additionally, we pointed out that there is also a risk on the other side (the opposite problem) of having too much dependency between a theory’s predictions and a scientist’s inferences, which would create theories of consciousness that are trivial and unfalsifiable. In other words, there are two “horns” to the problem of falsifiability for theories of consciousness.
@@
The dual-horned “Kleiner-Hoel dilemma” established in [45] goes like this: A successful theory of consciousness must not be a priori falsified by the class of arguments like the Unfolding Argument (broadly, any such “universal substitution,” like unfolding, that creates a pathological circumstance of changing all, or at least the vast majority, of predictions, while holding inferences constant). However, the same theory must also not be unfalsifiable in principle. This creates only a small space through which to pass; theories of consciousness are “caught between Scylla and Charybdis.”
@@
On one side, theories like IIT [68] or recurrent processing theories [56] are a priori falsified by substitutions like unfolded neural networks (since pretty much all the predictions of a theory would change under such a substitution, while the experimenter’s inferences based on reports or behavior would stay exactly the same, creating universal mismatches between predictions and inferences and thus resulting in falsification). Meanwhile, theories like behaviorism or global functionalism (both being based solely on input/output mattering for consciousness) are unfalsifiable, because predictions of the theories and inferences about consciousness from experiments can never differ (e.g., if your theory of consciousness is that consciousness is identical to behaving as if conscious, how could it ever be falsified?).
@@
This was a strong challenge to the consciousness literature. The primary reply against it has been simply that the Kleiner-Hoel dilemma is too strong and therefore cannot possibly be true, at least without ruling out consciousness research in general [72]. This is a debatable reason to reject a conclusion, but more importantly, it falls by the wayside once an appropriate theory is introduced (even a sketch of one) that can navigate the Kleiner-Hoel dilemma. Critically, if the Kleiner-Hoel dilemma is navigated successfully in humans, but provably cannot be navigated in LLMs (due to their close resemblance to systems that we know cannot be targets for non-trivial and falsifiable theories), this would form a strong argument against LLM consciousness. In other words, there may be properties of human brains that allow them to “pass through” the Kleiner-Hoel dilemma unscathed, but LLMs, lacking those properties, cannot.
That is what this paper shows. It is structured in three parts.
First, I overview a formal framework for testing theories of consciousness [45], and demonstrate how the Kleiner-Hoel dilemma naturally arises from it, including a new proof based on Kolmogorov’s notion of a “shortest program” for how even functionalist theories face challenges from substitutions (or fall into triviality).
@@
Second, I show that the Kleiner-Hoel dilemma has extremely strong argumentative force in LLMs, because LLMs are much closer than human brains in “substitution space” to input/output equivalent substitutions like lookup tables that non-trivial theories of consciousness can’t possibly judge as conscious. That is, there is no theory of consciousness available for contemporary LLMs that would not fall on one of the horns of the Kleiner-Hoel dilemma. Therefore, there is no scientific theory of consciousness which could judge contemporary LLMs as conscious. Therefore, they are not conscious.
@@
The strength of this argument is that it is entirely agnostic to any assumptions about particular theories of consciousness being true, nor does it make assumptions about the “specialness” of biology, or the possibility of AI consciousness in general, and has few philosophical commitments (other than that theories of consciousness should be falsifiable and non-trivial).
@@
Third, this paper shows that the Kleiner-Hoel dilemma has a potential solution in humans by grounding a theory of consciousness not in causal structure like IIT, nor functional or dynamical features like Global Workspace Theory, but in continual learning. This fits well with the recent observation by O’Reilly-Shah et al. that a static FNN cannot act as a universal substitution for a plastic RNN [59]. However, those recent proofs only show that plastic RNNs are protected from the specifics of the Unfolding Argument—they do not fully distinguish whether the recurrence or plasticity is the cause, nor ask if this protection extends to other universal substitutions. Here I show that grounding a theory of consciousness in the broader context of continual learning explicitly avoids both horns of the Kleiner-Hoel dilemma.
@@
This result implies that scientific theories of consciousness are, until other cases that avoid the Kleiner-Hoel dilemma are identified, strongly constrained to not merely being confined to systems that can learn (like plastic RNNs), but more generally that we can constrain theories of consciousness to, at minimum, involve continual learning itself. Such a result significantly winnows the field of consciousness research and illuminates the path toward less popular, but ultimately more scientific, theories of consciousness. It also implies that the vast majority of the search for the neural correlates of consciousness may be incommensurate, because testing static input/output behavior alone is not how one empirically tests for consciousness.
@@
Intriguingly, this research indicates that one major reason deployed LLMs lack consciousness is their static structure, and that unlocking continual learning, already a significant target in AI research [53], could possibly change this disproof. It also provides grounds to believe that the current cognitive limitations of LLMs in adaptation to novelty and generalization are based on their lack of consciousness, which is connected to their lack of continual learning.
3 A Formal Framework for Falsifying Theories of Consciousness
@@
The line of reasoning here will make use of a previously developed framework for falsification of theories of consciousness [45]. This is a formalized mathematical framework for testing theories of consciousness, based on treating theories as mapping a space of possible experiences onto physical predictions. Its advantage is that it is abstract, general, and (relatively) simple.
@@
Here, I leave much of the background logic and reasoning to previous work and deal only with the high-level details that are relevant (see [45] for those details). For our purposes what matters is that testing a theory of consciousness is based on comparing two functions: predictions and inferences.
@@
First, what are predictions? Predictions pick out a particular element of the space of possible experiences, $E$. The correspondence (here, for simplicity, referred to as a function) that picks out experiences is a mapping from experimental data about the system’s workings, $O$ (such as neuroimaging data), to the space of experiences, $E$; this mapping is dubbed pred, for “prediction.”
@@
Theories are defined by their unique methods of making predictions. However, predictions alone cannot falsify a theory. They need to be compared to empirical inferences about consciousness, either explicit or implicit.
Usually, these inferences are based on verbal report, behavior, or, more recently, even “no-report paradigms” [70]. But even “no-report paradigm” cases (based in things like eye movement) rely on correlating with report first. In a psychology lab, button presses to indicate what was flashed in a masking stimulus are inference data; in another context, like in assessing LLM consciousness, the outputs of an artificial neural network could count as inference data. Inferences are, essentially, empirical reasons to believe a system might be conscious irrespective of any given theory or class of theories. Perhaps the clearest inference would be an explicit claim by the system to consciousness itself, but individual examples of inferences while testing a theory of consciousness might be, e.g., correctly reporting sensory inputs, responding to pain, etc.
Overall, inferences can be represented by another function, which maps between the data generated by observations about the behavior of a system (like its input/output) and the space of experiences $E$:
@@
Experimental DataExperiencesData fromverbal reportsData aboutsystem’s workingsEEinferencespredictions
Figure 1: A theory of consciousness is not falsified (sometimes called “experimentally confirmed,” although that language is often not appropriate) when inference from behavioral reports and prediction from neural data converge on the same experience from $E$.
Obviously, this formal setup is an extremely idealized and abstracted version of testing a theory of consciousness (for a defense of the irrelevancy of other details, however, see [45]).
It is important that a falsification framework not be overfitted or over-sensitive. For the purposes herein, I do not require that falsification triggers from just any possible mismatch between pred and inf. For instance, imagine a theory of consciousness that performs very well in humans and we are starting to believe scientifically. Then, we apply it to some systems for which we cannot access consciousness as well (say, plants). This gives us a (to many) surprising result: that plants are conscious to some degree. However, it may be that no experimental inference from plants can fully support the weight of the claim of the theory, and therefore, there are technically some mismatches. The point here is not to say this should happen with a theory, or is totally non-problematic, but just that, at this time, the falsification framework introduced herein can be optionally permissive about some occasional mismatches.
On the other hand, if, e.g., a theory of consciousness predicts that extremely simple systems have degrees of consciousness greater than that of a human [1, 69], such a radical mismatch could count as actual falsification—but only in light of the inference that a system’s behavioral or input/output simplicity is evidence against such a high degree of consciousness. Ideally, such inferences should always be made explicit when reasoning about consciousness through the lens of falsification. However, it is important to note that keeping assumptions about inferences as neutral as possible is important, or else testing a theory of consciousness would lose its desired empirical “universality” [42].
@@
Figure 2: The Substitution Argument: Changing the system (and thus significantly changing the prediction function), either by transformation or swapping in a different system, all while holding inference data constant, should always be possible as long as inferences and predictions are completely independent. Here, following a substitution, the predicted experience shifts from $E$ to $E^{\prime}$, but the inferred experience remains $E$, creating the mismatches that falsify the theory.
Importantly, it is obvious that if inf is held constant, but pred is varied significantly, then this will create consistent divergence (mismatches) across the entirety (or vast majority) of a theory’s predictions. This is pathological behavior and should necessarily falsify a given theory, as now every prediction is wrong (or the vast majority are wrong) for a system with identical (or, again, extremely similar) inferences. Specifically, in [45] we took the approach that a “universal substitution” (of one system for another) in an experimental setup (within sensible allowances), but in which inferences do not change—but all predictions do change—would completely falsify a particular theory. We called such a possibility a “universal substitution” and confined the analyses to such cases in order to avoid more debatable instances. E.g., it is unclear to what degree “partial substitutions” triggering mismatches should count as falsification, since it may make sense to sometimes err on the side of a well-confirmed theory in edge cases (but again, this question can be set aside here).
@@
As we showed, substitutions are guaranteed to exist as long as predictions and inferences are totally independent. Another way of saying this: Just from knowing the structure of Integrated Information Theory, and that the calculation of $\Phi$ is entirely screened off from the input/output of a system, there can be systems wherein $\Phi$ differs but the input/output does not. However, the reason why this appears to be such a problem for IIT is simply that IIT is formally well-developed. Other theories of consciousness will eventually become equally formally developed, and have to specify similarly precise predictions to $\Phi$, and then they will fall victim to the same problems.